# HiStream: Efficient High-Resolution Video Generation via Redundancy-Eliminated Streaming

### v1

@@ 1 Introduction
With the rapid development of video diffusion models Blattmann et al. (2023); Yang et al. (2024); Chen et al. (2024); HaCohen et al. (2024); Kong et al. (2024); Wan et al. (2025), high-resolution video generation is increasingly essential for applications in film production, virtual reality, and digital media, where visual realism and fine-grained details are critical. However, generating high-fidelity and temporally coherent videos remains extremely challenging due to the substantial spatio-temporal complexity. Although recent high-resolution generation methods Zhang et al. (2025a); Ren et al. (2025); Qiu et al. (2024, 2025a) have demonstrated impressive success, their computational efficiency is far from practical, as the inference cost scales quadratically based on the video spatial resolution and temporal length. As a result, the challenge of generating realistic 1080p videos in a time-efficient manner remains unsolved.
To address these computational bottlenecks, several strategies have been proposed. Sampling acceleration techniques, such as timestep distillation Wang et al. (2023b); Sauer et al. (2024); Yin et al. (2024b, a), effectively reduce the number of denoising steps. However, the computational cost within each step remains high when applied to high-resolution videos. Concurrently, sparse or sliding attention mechanisms Zhang et al. (2025b); Cai et al. (2025); Huang et al. (2025) mitigate the full quadratic cost, but their cache and memory requirements still grow linearly with spatiotemporal volume. Thus, while these methods offer partial acceleration in low-resolution or short-video settings, they remain insufficient for scaling to long-duration, high-fidelity 1080p video generation.
We argue that the fundamental source of this inefficiency is inherent computational redundancy across spatial, temporal, and timestep axes. Spatially, significant computation is wasted in early denoising steps, which primarily establish coarse motion and structure Liu et al. (2025b) and can be simply approximated at a low-resolution. High-resolution details are only recovered in later steps. This insight directly motivates our Dual-Resolution Caching mechanism. Temporally, similar redundancy exists in autoregressive models. We find that only the initial frame (acting as a persistent keyframe, akin to an attention sink Gu et al. (2024)) and a small set of neighboring frames are crucial for temporal consistency. This observation leads to our Anchor-Guided Sliding Window strategy. Finally, we observe that the denoising trajectory differs significantly between the initial and subsequent chunks, a phenomenon that motivates our Asymmetric Denoising Strategy.
@@
Extensive experiments on 1080p video benchmarks demonstrate that our primary HiStream model (utilizing the first two optimizations) attains state-of-the-art visual quality with negligible quality loss, while offering up to $76.2\times$ faster denoising compared to the base video foundation model Wan2.1 Wan et al. (2025). Our faster variant, HiStream+, which additionally employs Asymmetric Denoising, pushes this acceleration to a remarkable $107.5\times$ over the baseline. This variant provides a powerful speed-quality trade-off, making HiStream a practical and scalable solution for both high-fidelity and ultra-fast high-resolution video generation.
@@ 2 Related Work
Diffusion-based methods have recently emerged as the leading approach for video generation. VDM Ho et al. (2022) first introduced diffusion models to this domain, establishing a foundation for subsequent research. Later works Blattmann et al. (2023); He et al. (2022); Guo et al. (2023); Chen et al. (2023); Wang et al. (2023a, c) built upon this framework with hierarchical latent designs, enabling the creation of longer videos with improved temporal consistency. However, these methods still rely on the UNet Ronneberger et al. (2015) architecture, which suffers from limited scalability. A recent shift towards Transformer-based architectures, specifically Diffusion Transformers (DiT) Peebles and Xie (2023), has emerged in video generation. Models like CogVideoX Yang et al. (2024) and Pyramid Flow Jin et al. (2024) were among the first to demonstrate the superior scalability and strong performance of this approach. Leveraging the scalability of DiT, models like LTX HaCohen et al. (2024), Hunyuan Kong et al. (2024), and Wan Wan et al. (2025) now generate visually realistic videos from text, surpassing prior UNet-based approaches.
@@
High resolution visual synthesis remains a long-standing challenge in the generative field due to the high-resolution data scarcity and the substantial computational cost of modeling, particularly for videos. Recent efforts have diverged into two main categories: (1) Training-based methods, which fine-tune on high-resolution data directly with large-size models Teng et al. (2023); Hoogeboom et al. (2023); Ren et al. (2024); Liu et al. (2024a); Guo et al. (2024); Zheng et al. (2024); Cheng et al. (2024); Zhang et al. (2025a); Ren et al. (2025); and (2) Tuning-free methods, which modify architecture to improve resolution without additional training data Haji-Ali et al. (2024); Lin et al. (2024a); Lee et al. (2023); Jin et al. (2023); Zhang et al. (2024); Kim et al. (2024). While directly training on high-resolution data is a fundamental solution, the scarcity of such data and the immense memory requirements make it impractical. Consequently, tuning-free methods have also become a practical choice.
However, most tuning-free approaches He et al. (2024); Du et al. (2024a); Huang et al. (2024a); Lin et al. (2024b); Liu et al. (2024b); Qiu et al. (2024) are designed for UNet-based models, focusing on mitigating receptive-field limitations that cause artifacts. DiT-based models face a distinct challenge, where misaligned positional encodings can cause blurriness at high resolutions. As foundational models shift to DiT architectures, tuning-free methods Du et al. (2024b); Qiu et al. (2025a) for them are crucial. A common alternative to native synthesis is a two-stage approach using external super-resolution models Wang et al. (2021); Zhou et al. (2024); Wang et al. (2025). While computationally cheap, this pipeline often fails to render fine details, yielding less faithful visual quality.
@@
The slow inference speed of iterative diffusion models has spurred multiple efficiency efforts. The most direct approach is reducing denoising steps, evolving from simple caching of redundant timesteps Lv et al. (2024); Liu et al. (2025a, c) to more powerful distillation techniques Luo et al. (2023); Wang et al. (2023b); Sauer et al. (2024); Yin et al. (2024b, a). Concurrently, other works have focused on reducing intra-network redundancy via sparsification Zhang and Agrawala (2025); Zhang et al. (2025b); Cai et al. (2025). A more recent and relevant trend is the acceleration of temporal self-attention. By reformulating video diffusion models causally Teng et al. (2025); Chen et al. (2025); Yin et al. (2025); Huang et al. (2025), these methods enable incremental inference using KV caching. This development mirrors discoveries in LLMs, namely the “attention sink” phenomenon Xiao et al. (2023); Gu et al. (2024); Qiu et al. (2025b). This property was famously exploited by StreamingLLM Xiao et al. (2023) to implement a local sliding-window attention with a persistent anchor token, achieving stable inference with a fixed-size KV cache. HiStream builds directly on this lineage of ideas, adapting them to the unique challenges of high-resolution video.
@@ 3 Methodology
(a) Autoregressive Inference with Anchor-Guided Sliding Window
@@ 3.1 Preliminaries
Our HiStream framework is an optimization of an autoregressive video diffusion model. This baseline architecture models the joint distribution of $N$ video frames $\mathbf{x}^{1:N}$ by factorizing it autoregressively: $p(\mathbf{x}^{1:N})=\prod_{i=1}^{N}p(\mathbf{x}^{i}\mid\mathbf{x}^{<i})$. Generation is a sequential process where each frame $\mathbf{x}^{i}$ is denoised from Gaussian noise, conditioned on the preceding frames $\mathbf{x}^{<i}$. In this work, our framework generates video in chunks of frames, rather than processing single frames individually. Our baseline WAN-2.1 is a Diffusion Transformer (DiT) Peebles and Xie (2023) that operates in the latent space of a causal 3D VAE Wan et al. (2025). The autoregressive factorization is implemented efficiently within the Transformer using causal attention masks, which restrict the model to only attend to past and current frames. This model is trained using the standard frame-wise diffusion loss, minimizing the MSE between the predicted noise $\hat{\mathbf{\epsilon}}_{\theta}$ and the true sampled noise $\mathbf{\epsilon}$:
@@
A key strategy for accelerating diffusion models is consistency distillation, which trains a student model to produce high-quality samples in one or a few steps, mimicking a multi-step teacher. While traditional consistency distillation can involve complex multi-step losses, Flow Matching (FM) Lipman et al. (2022) provides a more direct and stable training objective. FM trains the student model $f_{\phi}$ to learn a conditional vector field that directly maps a noisy input $\mathbf{x}_{t}$ to the clean data $\mathbf{x}_{0}$. This simplifies the objective to a stable $L_{2}$ loss, making it a robust method for few-step inference with the conditioned information $\mathbf{c}$:
@@ 3.2 Spatial Compression
Based on this insight, our DRC strategy re-architects the denoising trajectory. Steps 1-2 are performed at a low resolution to efficiently establish the global structure. The latent representation is then upsampled to the target high resolution, and Steps 3-4 perform high-resolution refinement. To enable a single network to handle this resolution shift, we adapt techniques validated in LLMs Qiu et al. (2025a), integrating Rotary Positional Embeddings (RoPE) with NTK-inspired scaling Peng et al. (2023) for generalizing positional information across scales. This progressive upscaling approach reserves the most expensive high-resolution operations for the final, critical refinement stages.
@@ 3.3 Temporal Compression
To ensure long-range consistency, we leverage the “attention sink” phenomenon observed in LLMs and autoregressive generation. Efficiency-focused models like StreamingLLM Xiao et al. (2023) strategically leverage this by using fixed anchor tokens to stabilize a sliding window. We observe that this phenomenon also naturally emerges in autoregressive video generation, with attention strongly concentrating on the first frame. This insight is so fundamental that, even without fine-tuning, dropping intermediate, low-value frames still yields high-quality results closely approximating full-history generation (Figure 4).
@@ 4 Experiments
Wan2.1 Wan et al. (2025) 1.3B 81 832×480 84.13 80.05 83.32 2.04
Self Forcing Huang et al. (2025) 1.3B 81 832×480 84.85 80.91 84.06 0.10
@@
Wan2.1 Wan et al. (2025) 1.3B 81 1920×1088 81.74 76.39 80.67 36.56
Self Forcing Huang et al. (2025) 1.3B 81 1920×1088 84.65 79.97 83.71 1.18
LTX HaCohen et al. (2024) 2B 81 1920×1088 80.31 69.54 78.16 1.60
FlashVideo Zhang et al. (2025a) 5B 49 1920×1072 83.75 81.62 83.32 6.40
@@
Figure 5: Qualitative comparisons. We visualize high-resolution videos generated by HiStream (Ours) against those by Wan2.1 Wan et al. (2025), Self Forcing Huang et al. (2025), LTX HaCohen et al. (2024), and FlashVideo Zhang et al. (2025a). The videos generated by HiStream exhibit the highest visual fidelity and the cleanest texture, free from spurious patterns or visible artifacts. Best viewed ZOOMED-IN.
@@
We design HiStream on top of the Wan2.1-T2V-1.3B model Wan et al. (2025), whose original resolution is $832\times 480$. Following the Self Forcing paradigm Huang et al. (2025), we leverage the pre-trained Wan2.1-T2V-14B model as a teacher to guide our student model’s training. Notably, HiStream requires no additional real video during this process. All training prompts are obtained from VidProM Wang and Yang (2024) as we follow Self Forcing Huang et al. (2025). We tune the model at $960\times 544$ resolution (memory limitation), while generating content at $1920\times 1088$ during inference. HiStream uses $4$-step diffusion and implements chunk-wise autoregressive variants, generating chunks of $M=3$ latent frames.
@@
We perform extensive evaluations using VBench Huang et al. (2024b) and a user preference study, covering both visual fidelity and semantic alignment. We use augmented prompts from VBench as all baselines require. We also rigorously examine the runtime efficiency of our approach to validate its applicability.
@@ 4.1 Comparison with Existing Baselines
We compare our model with relevant open-source efficient/high-resolution video generation models of similar scale. Our comparisons include three diffusion models: Wan2.1-1.3B Wan et al. (2025) (our base model structure), Self Forcing Huang et al. (2025) (our initialization weights), LTX HaCohen et al. (2024) (known for efficiency), and FlashVideo Zhang et al. (2025a) (known for high-resolution generation). Table 4 exhibits that our method achieves the best Quality Score and Total Score on VBench Huang et al. (2024b) in the high-resolution generation. While scores in high-resolution generation are generally lower than those at low resolutions, our method nonetheless achieves a score that is highly competitive, approaching the best result recorded in the low-resolution setting. In terms of Per-Frame Denoising Latency, our proposed HiStream is the most efficient, achieving a minimal latency of only $\mathbf{0.48\,s}$. This performance represents an impressive $\mathbf{76.2\times}$ acceleration compared to the baseline Wan2.1-T2V-1.3B (36.56 s). Crucially, HiStream also significantly outperforms the existing efficient method Self Forcing (1.18 s), demonstrating a near $\mathbf{2.5\times}$ speedup. As shown in Figure 5, the videos generated by HiStream exhibit the highest visual fidelity and the cleanest texture, free from spurious patterns or visible artifacts.

### v2

@@ 1 Introduction
With the rapid development of video diffusion models (Blattmann et al., 2023; Yang et al., 2024; Chen et al., 2024; HaCohen et al., 2024; Kong et al., 2024; Wan et al., 2025), high-resolution video generation is increasingly essential for applications in film production, virtual reality, and digital media, where visual realism and fine-grained details are critical. However, generating high-fidelity and temporally coherent videos remains extremely challenging due to the substantial spatio-temporal complexity. Although recent high-resolution generation methods (Zhang et al., 2025a; Ren et al., 2025; Qiu et al., 2024, 2025a) have demonstrated impressive success, their computational efficiency is far from practical, as the inference cost scales quadratically based on the video spatial resolution and temporal length. As a result, the challenge of generating realistic 1080p videos in a time-efficient manner remains unsolved.
To address these computational bottlenecks, several strategies have been proposed. Sampling acceleration techniques, such as timestep distillation (Wang et al., 2023b; Sauer et al., 2024; Yin et al., 2024b, a), effectively reduce the number of denoising steps. However, the computational cost within each step remains high when applied to high-resolution videos. Concurrently, sparse or sliding attention mechanisms (Zhang et al., 2025b; Cai et al., 2025; Huang et al., 2025) mitigate the full quadratic cost, but their cache and memory requirements still grow linearly with spatiotemporal volume. Thus, while these methods offer partial acceleration in low-resolution or short-video settings, they remain insufficient for scaling to long-duration, high-fidelity 1080p video generation.
We argue that the fundamental source of this inefficiency is inherent computational redundancy across spatial, temporal, and timestep axes. Spatially, significant computation is wasted in early denoising steps, which primarily establish coarse motion and structure (Liu et al., 2025b) and can be simply approximated at a low-resolution. High-resolution details are only recovered in later steps. This insight directly motivates our Dual-Resolution Caching mechanism. Temporally, similar redundancy exists in autoregressive models. We find that only the initial frame (acting as a persistent keyframe, akin to an attention sink (Gu et al., 2024)) and a small set of neighboring frames are crucial for temporal consistency. This observation leads to our Anchor-Guided Sliding Window strategy. Finally, we observe that the denoising trajectory differs significantly between the initial and subsequent chunks, a phenomenon that motivates our Asymmetric Denoising Strategy.
@@
Extensive experiments on 1080p video benchmarks demonstrate that our primary HiStream model (utilizing the first two optimizations) attains state-of-the-art visual quality with negligible quality loss, while offering up to $76.2\times$ faster denoising compared to the base video foundation model Wan2.1 (Wan et al., 2025). Our faster variant, HiStream+, which additionally employs Asymmetric Denoising, pushes this acceleration to a remarkable $107.5\times$ over the baseline. This variant provides a powerful speed-quality trade-off, making HiStream a practical and scalable solution for both high-fidelity and ultra-fast high-resolution video generation.
@@ 2 Related Work
Diffusion-based methods have recently emerged as the leading approach for video generation. VDM (Ho et al., 2022) first introduced diffusion models to this domain, establishing a foundation for subsequent research. Later works (Blattmann et al., 2023; He et al., 2022; Guo et al., 2023; Chen et al., 2023; Wang et al., 2023a, c) built upon this framework with hierarchical latent designs, enabling the creation of longer videos with improved temporal consistency. However, these methods still rely on the UNet (Ronneberger et al., 2015) architecture, which suffers from limited scalability. A recent shift towards Transformer-based architectures, specifically Diffusion Transformers (DiT) (Peebles and Xie, 2023), has emerged in video generation. Models like CogVideoX (Yang et al., 2024) and Pyramid Flow (Jin et al., 2024) were among the first to demonstrate the superior scalability and strong performance of this approach. Leveraging the scalability of DiT, models like LTX (HaCohen et al., 2024), Hunyuan (Kong et al., 2024), and Wan (Wan et al., 2025) now generate visually realistic videos from text, surpassing prior UNet-based approaches.
@@
High resolution visual synthesis remains a long-standing challenge in the generative field due to the high-resolution data scarcity and the substantial computational cost of modeling, particularly for videos. Recent efforts have diverged into two main categories: (1) Training-based methods, which fine-tune on high-resolution data directly with large-size models (Teng et al., 2023; Hoogeboom et al., 2023; Ren et al., 2024; Liu et al., 2024a; Guo et al., 2024; Zheng et al., 2024; Cheng et al., 2024; Zhang et al., 2025a; Ren et al., 2025); and (2) Tuning-free methods, which modify architecture to improve resolution without additional training data (Haji-Ali et al., 2024; Lin et al., 2024a; Lee et al., 2023; Jin et al., 2023; Zhang et al., 2024; Kim et al., 2024). While directly training on high-resolution data is a fundamental solution, the scarcity of such data and the immense memory requirements make it impractical. Consequently, tuning-free methods have also become a practical choice.
However, most tuning-free approaches (He et al., 2024; Du et al., 2024a; Huang et al., 2024a; Lin et al., 2024b; Liu et al., 2024b; Qiu et al., 2024) are designed for UNet-based models, focusing on mitigating receptive-field limitations that cause artifacts. DiT-based models face a distinct challenge, where misaligned positional encodings can cause blurriness at high resolutions. As foundational models shift to DiT architectures, tuning-free methods (Du et al., 2024b; Qiu et al., 2025a) for them are crucial. A common alternative to native synthesis is a two-stage approach using external super-resolution models (Wang et al., 2021; Zhou et al., 2024; Wang et al., 2025). While computationally cheap, this pipeline often fails to render fine details, yielding less faithful visual quality.
@@
The slow inference speed of iterative diffusion models has spurred multiple efficiency efforts. The most direct approach is reducing denoising steps, evolving from simple caching of redundant timesteps (Lv et al., 2024; Liu et al., 2025a, c) to more powerful distillation techniques (Luo et al., 2023; Wang et al., 2023b; Sauer et al., 2024; Yin et al., 2024b, a). Concurrently, other works have focused on reducing intra-network redundancy via sparsification (Zhang and Agrawala, 2025; Zhang et al., 2025b; Cai et al., 2025). A more recent and relevant trend is the acceleration of temporal self-attention. By reformulating video diffusion models causally (Teng et al., 2025; Chen et al., 2025; Yin et al., 2025; Huang et al., 2025), these methods enable incremental inference using KV caching. This development mirrors discoveries in LLMs, namely the “attention sink” phenomenon (Xiao et al., 2023; Gu et al., 2024; Qiu et al., 2025b). This property was famously exploited by StreamingLLM (Xiao et al., 2023) to implement a local sliding-window attention with a persistent anchor token, achieving stable inference with a fixed-size KV cache. HiStream builds directly on this lineage of ideas, adapting them to the unique challenges of high-resolution video.
@@ 3 Methodology
(a) AR Inference with Anchor-Guided Sliding Window
@@ 3.1 Preliminaries
Our HiStream framework is an optimization of an autoregressive video diffusion model. This baseline architecture models the joint distribution of $N$ video frames $\mathbf{x}^{1:N}$ by factorizing it autoregressively: $p(\mathbf{x}^{1:N})=\prod_{i=1}^{N}p(\mathbf{x}^{i}\mid\mathbf{x}^{<i})$. Generation is a sequential process where each frame $\mathbf{x}^{i}$ is denoised from Gaussian noise, conditioned on the preceding frames $\mathbf{x}^{<i}$. In this work, our framework generates video in chunks of frames, rather than processing single frames individually. Our baseline WAN-2.1 is a Diffusion Transformer (DiT) (Peebles and Xie, 2023) that operates in the latent space of a causal 3D VAE (Wan et al., 2025). The autoregressive factorization is implemented efficiently within the Transformer using causal attention masks, which restrict the model to only attend to past and current frames. This model is trained using the standard frame-wise diffusion loss, minimizing the MSE between the predicted noise $\hat{\mathbf{\epsilon}}_{\theta}$ and the true sampled noise $\mathbf{\epsilon}$:
@@
A key strategy for accelerating diffusion models is consistency distillation, which trains a student model to produce high-quality samples in one or a few steps, mimicking a multi-step teacher. While traditional consistency distillation can involve complex multi-step losses, Flow Matching (FM) (Lipman et al., 2022) provides a more direct and stable training objective. FM trains the student model $f_{\phi}$ to learn a conditional vector field that directly maps a noisy input $\mathbf{x}_{t}$ to the clean data $\mathbf{x}_{0}$. This simplifies the objective to a stable $L_{2}$ loss, making it a robust method for few-step inference with the conditioned information $\mathbf{c}$:
@@ 3.2 Spatial Compression
Based on this insight, our DRC strategy re-architects the denoising trajectory. Steps 1-2 are performed at a low resolution to efficiently establish the global structure. The latent representation is then upsampled to the target high resolution, and Steps 3-4 perform high-resolution refinement. To enable a single network to handle this resolution shift, we adapt techniques validated in LLMs (Qiu et al., 2025a), integrating Rotary Positional Embeddings (RoPE) with NTK-inspired scaling (Peng et al., 2023) for generalizing positional information across scales. This progressive upscaling approach reserves the most expensive high-resolution operations for the final, critical refinement stages.
@@ 3.3 Temporal Compression
To ensure long-range consistency, we leverage the “attention sink” phenomenon observed in LLMs and autoregressive generation. Efficiency-focused models like StreamingLLM (Xiao et al., 2023) strategically leverage this by using fixed anchor tokens to stabilize a sliding window. We observe that this phenomenon also naturally emerges in autoregressive video generation, with attention strongly concentrating on the first frame. This insight is so fundamental that, even without fine-tuning, dropping intermediate, low-value frames still yields high-quality results closely approximating full-history generation (Figure 4).
@@ 4 Experiments
Wan2.1 (Wan et al., 2025) 1.3B 81 832×480 84.13 80.05 83.32 2.04
Self Forcing (Huang et al., 2025) 1.3B 81 832×480 84.85 80.91 84.06 0.10
@@
Wan2.1 (Wan et al., 2025) 1.3B 81 1920×1088 81.74 76.39 80.67 36.56
Self Forcing (Huang et al., 2025) 1.3B 81 1920×1088 84.65 79.97 83.71 1.18
LTX (HaCohen et al., 2024) 2B 81 1920×1088 80.31 69.54 78.16 1.60
FlashVideo (Zhang et al., 2025a) 5B 49 1920×1072 83.75 81.62 83.32 6.40
@@
Figure 5: Qualitative comparisons. We visualize high-resolution videos generated by HiStream (Ours) against those by Wan2.1 (Wan et al., 2025), Self Forcing (Huang et al., 2025), LTX (HaCohen et al., 2024), and FlashVideo (Zhang et al., 2025a). The videos generated by HiStream exhibit the highest visual fidelity and the cleanest texture, free from spurious patterns or visible artifacts. Best viewed ZOOMED-IN.
@@
We design HiStream on top of the Wan2.1-T2V-1.3B model (Wan et al., 2025), whose original resolution is $832\times 480$. Following the Self Forcing paradigm (Huang et al., 2025), we leverage the pre-trained Wan2.1-T2V-14B model as a teacher to guide our student model’s training. Notably, HiStream requires no additional real video during this process. All training prompts are obtained from VidProM (Wang and Yang, 2024) as we follow Self Forcing (Huang et al., 2025). We tune the model at $960\times 544$ resolution (memory limitation), while generating content at $1920\times 1088$ during inference. HiStream uses $4$-step diffusion and implements chunk-wise autoregressive variants, generating chunks of $M=3$ latent frames.
@@
We perform extensive evaluations using VBench (Huang et al., 2024b) and a user preference study, covering both visual fidelity and semantic alignment. We use augmented prompts from VBench as all baselines require. We also rigorously examine the runtime efficiency of our approach to validate its applicability.
@@ 4.1 Comparison with Existing Baselines
We compare our model with relevant open-source efficient/high-resolution video generation models of similar scale. Our comparisons include three diffusion models: Wan2.1-1.3B (Wan et al., 2025) (our base model structure), Self Forcing (Huang et al., 2025) (our initialization weights), LTX (HaCohen et al., 2024) (known for efficiency), and FlashVideo (Zhang et al., 2025a) (known for high-resolution generation). Table 4 exhibits that our method achieves the best Quality Score and Total Score on VBench (Huang et al., 2024b) in the high-resolution generation. While scores in high-resolution generation are generally lower than those at low resolutions, our method nonetheless achieves a score that is highly competitive, approaching the best result recorded in the low-resolution setting. In terms of Per-Frame Denoising Latency, our proposed HiStream is the most efficient, achieving a minimal latency of only $\mathbf{0.48\,s}$. This performance represents an impressive $\mathbf{76.2\times}$ acceleration compared to the baseline Wan2.1-T2V-1.3B (36.56 s). Crucially, HiStream also significantly outperforms the existing efficient method Self Forcing (1.18 s), demonstrating a near $\mathbf{2.5\times}$ speedup. As shown in Figure 5, the videos generated by HiStream exhibit the highest visual fidelity and the cleanest texture, free from spurious patterns or visible artifacts.